# What Moves Together Belongs Together

### v1

@@ Abstract
00footnotetext: $*{}^{*}$ Correspondence to j.seidenschwarz@tum.de.
@@ 2 Related Work
Figure 2: Segment Moving in Lidar for Pseudo-Labeling: We first preprocess the point cloud to remove moving points and predict per-point trajectories on the filtered point cloud (preprocessing and trajectory prediction). Then, we extract velocity-based features from the trajectories and learn to cluster, i.e., segment points based on motion-patters using a Message-Passing Netowrk [12] in a fully data-driven manner (SeMoLi). From point segments, we extract bounding boxes and inflate them (extracting and inflating bounding boxes). Finally, we apply our approach on unlabeled Lidar streams to obtain pseudo-labels, that we use to train object detectors.
@@ 4.2 Ablations Studies
Graph
Oracle Velocity kNN 35.7 67.1 46.6 39.4 74.1 51.4
@@
Table 3: Pseudo-label quality comparison (3DIoU): We compare our SeMoLi to different variants of DBSCAN [30], augmented with scene flow (DBSCAN++), long-term trajectory information (DBSCAN++${}_{l}$) and outlier filtering (normal-$\dagger$). In gray (top) we report results using ground truth scene flow and trajectories, and below we report scene flow and motion trajectories obtained using [43]. SeMoLi consistently performs favorably compared to all DBSCAN variants, when using perfect “oracle” motion cues, as well as when using the estimated (noisy) scene flow method.
@@ 4.3 Evaluation of Pseudo-Label Quality
Baselines.  We compare SeMoLi to three baselines: vanilla DBSCAN, DBSCAN++ (cf., [30]), and its variant DBSCAN++${}_{l}$ that utilizes our (long-term) velocity-based motion feature for a fair comparison. ${}^{\dagger}$ indicates heuristic filtering based on bounding box dimensions (detailed in the appendix). We also report variant of our method that utilizes local motion cues (scene flow), SeMoLi ${}_{\texttt{sf}}$.
${}^{\dagger}$ SeMoLi ${}^{\texttt{10}}$
@@
${}^{\dagger}$ SeMoLi ${}^{\texttt{90}}$
@@
All (Mov. + stat.)
Stat. + Mov., class-specific 0 100 35.5 55.5 37.1 69.7 44.6 80.3
@@
Moving only
Stat. + Mov., class-specific 0 100 30.5 34.5 43.2 36.4 41.1 85.6
@@ 9 Registration
For a fair comparison with our baseline ${}^{\dagger}$ [30], we also implemented the registration [13] algorithm provided in the paper. The authors state that they track bounding boxes with a BEV IoU threshold of $0.1$, i.e., a small overlap is enough for two bounding boxes to be matched together. Then they register bounding boxes over all tracks to obtain a more complete representation of the underlying object. Contrary to the performance reported in [30], adding registration to our pipeline worsened our performance (see Tab. 11). We assume that this is due to our more noisy point cloud filtering. As soon as one detection in a track is noisy, due to the registration all bounding boxes in a track will be noisy. We also show the performance of our best registration setting, where we: (i) utilize a threshold of $0.5$, (ii) only apply constrained ICP to obtain transformations between detections but propagate the dimensions of the most dense point cloud, (iii) only apply constrained ICP on clusters with more than $50$ points and tracks with a length $\geq 5$. However this also does not improve our performance. Hence, we rely on bounding box inflation as our only post-processing step due to its simplicity and high efectiveness.
@@ 10 Details DBSCAN and DBSCAN++ re-implementation
Table 12: Comparison of ${}^{\dagger}$ with and without minimum number of points per cluster: We compare the performance of the re-implementation of our baseline ${}^{\dagger}$ as described in [30] where clusters are allowed to contain only one point. In our adapted re-implementation we define the minimum number of points per cluster for position- as well as scene flow-based clusterings as $10$ and the minimum number of samples of their intersection as $20$.
@@ 13 PP Performance with different percentages of pseudo and labeled data
In this section, we detail PointPillars (PP) performance on different percentages of labeled and pseudo-labeled data (see Tab. 13). For this we compare different versions of SeMoLi, i.e., SeMoLi ${}_{\texttt{10}}$, SeMoLi ${}_{\texttt{50}}$, and SeMoLi ${}_{\texttt{90}}$ with ${}^{\dagger}$ and our baselines trained on ground truth data.
@@
Training without Labeled Data.  We compare the performance when we train PP with pseudo-labels generated using ${}^{\dagger}$ and the different versions of SeMoLi on different amounts of pseudo-labeled data.
@@
Evaluation on static and moving objects.  Interestingly, for SeMoLi as well as for ${}^{\dagger}$ without labeled data, the best performance is achieved utilizing $50\%$ of the pseudo-labeled data. Precision as well as recall are always similar ${}^{\dagger}$ with a constantly low precision. This indicates that the noisy training signal leads to PP predicting a large number of false positive bounding boxes. Pseudo-labels generated by SeMoLi lead to increased precision as we increase the amount of pseudo-labeled data mirroring the high pseudo-label quality. However, recall drops which indicates that pseudo-labels generated by SeMoLi indeed focus on moving objects. The higher the ratio of moving objects PP obtains as supervisory signal the less static objects it will predict.
@@
Evaluation on moving objects.  On moving objects only, the performance of ${}^{\dagger}$ when utilizing only $10\%$ pseudo-labeled data is significantly worse than the performances utilizing $50\%,90\%$ or $100\%$ which are relatively stable. This indicates that ${}^{\dagger}$ pseudo-labels in general contain many noisy pseudo labels not belonging to moving objects and that the amount of moving objects is not enough in the $10\%$ split. Afterwards the ${}^{\dagger}$ pseudo-labels with low singal-to-noise ratio do not add to the training signal. For SeMoLi, we again observe a constant increase of precision the more pseudo-labeled data we utilize, indicating that the overall higher quality of the bounding boxes can be transferred to training PP.
@@
Training with Labeled Data.  Finally, we compare the performance when we train PP on a combination of labeled data and pseudo-labeled data generated using ${}^{\dagger}$ and the different versions of SeMoLi on different data splits.
@@
Evaluation on static and moving objects.  SeMoLi profits significantly more from adding labeled data to PP training with respect to precision than ${}^{\dagger}$. This indicates that the pseudo-labels generated by ${}^{\dagger}$ hamper the prediction of precise pseudo labels compared to when using SeMoLi’s pseudo-labels. This hypothesis is supported by the fact that the precision using ${}^{\dagger}$ does not significantly change depending on the amount of labeled data, while SeMoLi shows the highest precision when using the largest amount of labeled data. Since ${}^{\dagger}$ generates noisy pseudo-labels with many predictions not belonging to moving objects, the recall is highly similar to the recall when utilizing SeMoLi. This leads to an overall not significantly worse AP value of ${}^{\dagger}$.
@@
Evaluation on moving objects.  Evaluating on moving objects only leads to similar observations, except that the recall when utilizing SeMoLi’s pseudo-labels is higher compared to when utilizing ${}^{\dagger}$ pseudo-labels. This again indicates that indeed SeMoLi’s pseudo-labels focus on moving objects while ${}^{\dagger}$’s pseudo-labels also contain many pseudo-labels not belonging to moving objects – the vast majority belonging to noise.
@@
All (Mov. + stat.)
Ground Truth Baselines
@@
${}^{\dagger}$ 10 0 0.6 3.0 0.8 5.1 26.0 14.2
${}^{\dagger}$ 50 0 0.7 3.0 0.7 5.8 25.9 15.3
${}^{\dagger}$ 90 0 0.7 3.2 1.2 5.8 25.2 14.7
${}^{\dagger}$ 100 0 0.8 3.4 0.8 5.9 25.9 14.9
@@
${}^{\dagger}$ 10 90 7.2 35.2 31.3 11.4 55.2 52.2
${}^{\dagger}$ 50 50 6.9 35.1 31.2 10.9 54.9 51.7
${}^{\dagger}$ 90 10 7.4 35.1 31.1 11.5 55.0 52.0
@@
Moving only
Ground Truth Baselines
@@
${}^{\dagger}$ 10 0 0.5 8.7 2.4 3.0 56.4 39.7
${}^{\dagger}$ 50 0 0.6 9.2 2.1 3.5 58.7 44.2
${}^{\dagger}$ 90 0 0.6 9.5 1.1 3.6 58.4 43.5
${}^{\dagger}$ 100 0 5.9 9.7 2.4 3.6 58.6 43.2
@@
${}^{\dagger}$ 10 90 2.0 34.5 30.1 4.0 70.5 61.6
${}^{\dagger}$ 50 50 1.9 34.3 29.6 3.8 70.2 60.2
${}^{\dagger}$ 90 10 2.0 34.5 29.8 4.1 70.1 61.0
@@
Table 13: Semi-supervised 3D object detection on Waymo Open Dataset: We evaluate models on all (top) and only moving (bottom) on Waymo Open validation set. For each evaluation procedure we (i) show the ground truth baselines. Then we (ii) compare the different versions of SeMoLi utilizing the same train_detector training split. To compare SeMoLi to our baseline ${}^{\dagger}$ we (iii) show a comparison of both utilizing different percentages of pseudo-labeled data. Finally, we (iv) also show a comparison utilizing different combinations of pseudo-labeled and labeled data. % GT indicates the amount of labeled training data, % Pseudo indicates the amount of pseudo-labeled data.

### v2

# SeMoLi: What Moves Together Belongs Together

@@ Abstract
00footnotetext: Correspondence to j.seidenschwarz@tum.de.
@@ 2 Related Work
Figure 2: Segment Moving in Lidar for Pseudo-Labeling: We first preprocess the point cloud to remove static points and predict per-point trajectories on the filtered point cloud (preprocessing and trajectory prediction). Then, we extract velocity-based features from the trajectories and learn to cluster, i.e., segment points based on motion-patters using a Message-Passing Netowrk [12] in a fully data-driven manner (SeMoLi). From point segments, we extract bounding boxes and inflate them (extracting and inflating bounding boxes). Finally, we apply our approach on unlabeled Lidar streams to obtain pseudo-labels, that we use to train object detectors.
@@ 4.2 Ablations Studies
Graph Oracle Velocity kNN 35.7 67.1 46.6 39.4 74.1 51.4
@@
Table 3: Pseudo-label quality comparison (3DIoU): We compare our SeMoLi to different variants of DBSCAN [30], augmented with scene flow (DBSCAN++), long-term trajectory information (DBSCAN++l) and outlier filtering ($\dagger$). In gray (top) we report results using ground truth scene flow and trajectories, and below we report scene flow and motion trajectories obtained using [43]. SeMoLi consistently performs favorably compared to all DBSCAN variants, when using perfect “oracle” motion cues, as well as when using the estimated (noisy) scene flow method.
@@ 4.3 Evaluation of Pseudo-Label Quality
Baselines.  We compare SeMoLi to three baselines: vanilla DBSCAN, DBSCAN++ (cf., [30]), and its variant DBSCAN++l that utilizes our (long-term) velocity-based motion feature for a fair comparison. indicates heuristic filtering based on bounding box dimensions (detailed in the appendix). We also report variant of our method that utilizes local motion cues (scene flow), SeMoLi ${}_{\texttt{sf}}$.
DBSCAN++ SeMoLi ${}^{\texttt{10}}$
@@
DBSCAN++ SeMoLi ${}^{\texttt{90}}$
@@
All (Mov. + stat.) Stat. + Mov., class-specific 0 100 35.5 55.5 37.1 69.7 44.6 80.3
@@
Moving only Stat. + Mov., class-specific 0 100 30.5 34.5 43.2 36.4 41.1 85.6
@@ 9 Registration
For a fair comparison with our baseline DBSCAN++ [30], we also implemented the registration [13] algorithm provided in the paper. The authors state that they track bounding boxes with a BEV IoU threshold of $0.1$, i.e., a small overlap is enough for two bounding boxes to be matched together. Then they register bounding boxes over all tracks to obtain a more complete representation of the underlying object. Contrary to the performance reported in [30], adding registration to our pipeline worsened our performance (see Tab. 11). We assume that this is due to our more noisy point cloud filtering. As soon as one detection in a track is noisy, due to the registration all bounding boxes in a track will be noisy. We also show the performance of our best registration setting, where we: (i) utilize a threshold of $0.5$, (ii) only apply constrained ICP to obtain transformations between detections but propagate the dimensions of the most dense point cloud, (iii) only apply constrained ICP on clusters with more than $50$ points and tracks with a length $\geq 5$. However this also does not improve our performance. Hence, we rely on bounding box inflation as our only post-processing step due to its simplicity and high efectiveness.
@@ 10 Details DBSCAN and DBSCAN++ re-implementation
Table 12: Comparison of DBSCAN++ with and without minimum number of points per cluster: We compare the performance of the re-implementation of our baseline DBSCAN++ as described in [30] where clusters are allowed to contain only one point. In our adapted re-implementation we define the minimum number of points per cluster for position- as well as scene flow-based clusterings as $10$ and the minimum number of samples of their intersection as $20$.
@@ 13 PP Performance with different percentages of pseudo and labeled data
In this section, we detail PointPillars (PP) performance on different percentages of labeled and pseudo-labeled data (see Tab. 13). For this we compare different versions of SeMoLi, i.e., SeMoLi ${}_{\texttt{10}}$, SeMoLi ${}_{\texttt{50}}$, and SeMoLi ${}_{\texttt{90}}$ with DBSCAN++ and our baselines trained on ground truth data.
@@
Training without Labeled Data.  We compare the performance when we train PP with pseudo-labels generated using DBSCAN++ and the different versions of SeMoLi on different amounts of pseudo-labeled data.
@@
Evaluation on static and moving objects.  Interestingly, for SeMoLi as well as for DBSCAN++ without labeled data, the best performance is achieved utilizing $50\%$ of the pseudo-labeled data. Precision as well as recall are always similar forDBSCAN++ with a constantly low precision. This indicates that the noisy training signal leads to PP predicting a large number of false positive bounding boxes. Pseudo-labels generated by SeMoLi lead to increased precision as we increase the amount of pseudo-labeled data mirroring the high pseudo-label quality. However, recall drops which indicates that pseudo-labels generated by SeMoLi indeed focus on moving objects. The higher the ratio of moving objects PP obtains as supervisory signal the less static objects it will predict.
@@
Evaluation on moving objects.  On moving objects only, the performance of DBSCAN++ when utilizing only $10\%$ pseudo-labeled data is significantly worse than the performances utilizing $50\%,90\%$ or $100\%$ which are relatively stable. This indicates that DBSCAN++ pseudo-labels in general contain many noisy pseudo labels not belonging to moving objects and that the amount of moving objects is not enough in the $10\%$ split. Afterwards the DBSCAN++ pseudo-labels with low singal-to-noise ratio do not add to the training signal. For SeMoLi, we again observe a constant increase of precision the more pseudo-labeled data we utilize, indicating that the overall higher quality of the bounding boxes can be transferred to training PP.
@@
Training with Labeled Data.  Finally, we compare the performance when we train PP on a combination of labeled data and pseudo-labeled data generated using DBSCAN++ and the different versions of SeMoLi on different data splits.
@@
Evaluation on static and moving objects.  SeMoLi profits significantly more from adding labeled data to PP training with respect to precision than DBSCAN++. This indicates that the pseudo-labels generated by DBSCAN++ hamper the prediction of precise pseudo labels compared to when using SeMoLi’s pseudo-labels. This hypothesis is supported by the fact that the precision using DBSCAN++ does not significantly change depending on the amount of labeled data, while SeMoLi shows the highest precision when using the largest amount of labeled data. Since DBSCAN++ generates noisy pseudo-labels with many predictions not belonging to moving objects, the recall is highly similar to the recall when utilizing SeMoLi. This leads to an overall not significantly worse AP value of DBSCAN++.
@@
Evaluation on moving objects.  Evaluating on moving objects only leads to similar observations, except that the recall when utilizing SeMoLi’s pseudo-labels is higher compared to when utilizing DBSCAN++ pseudo-labels. This again indicates that indeed SeMoLi’s pseudo-labels focus on moving objects while DBSCAN++’s pseudo-labels also contain many pseudo-labels not belonging to moving objects – the vast majority belonging to noise.
@@
All (Mov. + stat.) Ground Truth Baselines
@@
DBSCAN++ 10 0 0.6 3.0 0.8 5.1 26.0 14.2
DBSCAN++ 50 0 0.7 3.0 0.7 5.8 25.9 15.3
DBSCAN++ 90 0 0.7 3.2 1.2 5.8 25.2 14.7
DBSCAN++ 100 0 0.8 3.4 0.8 5.9 25.9 14.9
@@
DBSCAN++ 10 90 7.2 35.2 31.3 11.4 55.2 52.2
DBSCAN++ 50 50 6.9 35.1 31.2 10.9 54.9 51.7
DBSCAN++ 90 10 7.4 35.1 31.1 11.5 55.0 52.0
@@
Moving only Ground Truth Baselines
@@
DBSCAN++ 10 0 0.5 8.7 2.4 3.0 56.4 39.7
DBSCAN++ 50 0 0.6 9.2 2.1 3.5 58.7 44.2
DBSCAN++ 90 0 0.6 9.5 1.1 3.6 58.4 43.5
DBSCAN++ 100 0 5.9 9.7 2.4 3.6 58.6 43.2
@@
DBSCAN++ 10 90 2.0 34.5 30.1 4.0 70.5 61.6
DBSCAN++ 50 50 1.9 34.3 29.6 3.8 70.2 60.2
DBSCAN++ 90 10 2.0 34.5 29.8 4.1 70.1 61.0
@@
Table 13: Semi-supervised 3D object detection on Waymo Open Dataset: We evaluate models on all (top) and only moving (bottom) on Waymo Open validation set. For each evaluation procedure we (i) show the ground truth baselines. Then we (ii) compare the different versions of SeMoLi utilizing the same train_detector training split. To compare SeMoLi to our baseline DBSCAN++ we (iii) show a comparison of both utilizing different percentages of pseudo-labeled data. Finally, we (iv) also show a comparison utilizing different combinations of pseudo-labeled and labeled data. % GT indicates the amount of labeled training data, % Pseudo indicates the amount of pseudo-labeled data.